# Convergence of Schrödinger operators on domains with scaled resonant potentials

### v1

@@ Abstract.
We consider Schrödinger operators on a bounded, smooth domain of dimension $d\geq 2$ with Dirichlet boundary conditions and a properly scaled potential, which depends only on the distance to the boundary of the domain. Our aim is to analyse the convergence of these operators as the scaling parameter tends to zero. If the scaled potential is resonant, the limit in strong resolvent sense is a Robin Laplacian with boundary coefficient expressed in terms of the mean curvature of the boundary. A counterexample shows that norm resolvent convergence cannot hold in general in this setting. If the scaled potential is non-negative (hence non-resonant), the limit in strong resolvent sense is the Dirichlet Laplacian. We conjecture that we can drop the non-negativity assumption in the non-resonant case.
@@ 1.1. Background and motivation
A typical example, where the limit depends on whether the scaled potential is resonant, is the approximation of a Schrödinger operator with a point $\delta$-interaction in three dimensions by a family of Schrödinger operators with scaled regular potentials. First partial results on this approximation are obtained in [AFH79, F72, N77]. To the best of our knowledge approximations of $\delta$-interactions were first addressed in full detail by Albeverio and Høegh-Krohn in [AH81], who proved the convergence in strong resolvent sense. The limit has a non-trivial point interaction if the scaled potential is resonant, otherwise the limit is the free Laplacian. By restricting to the class of radially symmetric potentials one obtains upon separation of variables a model problem on the half-line, which was considered separately by Šeba in [Š85]. He proved that a family of half-line Schrödinger operators with Dirichlet boundary conditions and locally scaled potentials converges in norm resolvent sense to the Neumann (or Robin) Laplacian on the half-line if the scaled potential is resonant and to the Dirichlet Laplacian otherwise. The main improvement in [Š85] is norm instead of strong resolvent convergence. The more general three-dimensional case was later also improved to norm resolvent convergence in the monograph [AGHH]. Further refinements and extensions of these results can be found in [DM16, SLS21].
@@
We remark that a similar phenomenon, where the limit qualitatively depends on whether the scaled potential is resonant or not, was observed in the approximation of $\delta^{\prime}$-potentials in a series of papers [GM09, GH13, G22], where the first two papers treat the one-dimensional case, while the last one deals with the two-dimensional case.
@@ 1.2. Resonant potentials in one dimension
Figure 1.1. The solutions $\psi_{0}$ of (1.1) for resonant and non-resonant potentials on the left and on the right, respectively.
@@
Our aim in the present paper is to address a multi-dimensional counterpart of these convergence results, which we will describe below in detail. We remark that the result of Šeba in [Š85] is more general than we stated here. He also shows how to approximate the one-dimensional Laplacian on the half-line with Robin boundary conditions by means of replacing $1/\varepsilon^{2}$ by $(1+\beta\varepsilon)/\varepsilon^{2}$ in the operator family ${\mathsf{H}}_{\varepsilon}$. We will not address this more general case in our analysis of the multi-dimensional problem, as we already obtain a Robin-type boundary conditions by geometry.
@@ 1.3. Main results
is called the shape operator; here $T_{s}\Sigma$ is the tangent space for $\Sigma$ at $s\in\Sigma$. The eigenvalues $\kappa_{1}(s),\kappa_{2}(s),\dots,\kappa_{d-1}(s)\colon\partial\Omega\to{\mathbb{R}}$ of ${\mathsf{L}}_{s}$ are called the principal curvatures of $\partial\Omega$. In our convention, all the principal curvatures are non-negative if and only if $\Omega$ is convex. The mean curvature of $\Sigma$ at $s\in\Sigma$ is defined as usual by
@@
We adopt the notation $H^{k}(\Omega)$ for the $L_{2}$-based Sobolev space on $\Omega$ of order $k\in{\mathbb{N}}$ and $H^{s}(\partial\Omega)$ for the $L_{2}$-based Sobolev space on the boundary $\partial\Omega$ of order $s\in{\mathbb{R}}_{+}$. For a function $u\in H^{2}(\Omega)$, we use the notation $u{\restriction}_{\Sigma}\in H^{3/2}(\Sigma)$ for its trace on the boundary and $\partial_{\nu}u{\restriction}_{\Sigma}\in H^{1/2}(\Sigma)$ for the normal derivative corresponding to the normal pointing outwards of $\Omega$. The role of the one-dimensional Neumann Laplacian ${\mathsf{H}}$ from the previous subsection is played now by the self-adjoint Robin Laplacian in $L_{2}(\Omega)$
@@
where the Robin coefficient is expressed in terms of the mean curvature. At the points, where the mean curvature vanishes, we recover locally Neumann boundary conditions. The role of the family of one-dimensional Schrödinger operators ${\mathsf{H}}_{\varepsilon}$ is played by the self-adjoint Schrödinger operator in $L_{2}(\Omega)$ defined for $\varepsilon>0$ by
@@
Assume that the potential $V\in C^{\infty}_{\mathrm{c}}(\overline{{{\mathbb{R}}}}_{+})$ is resonant in the sense of Definition 1.1. Then the family of scaled Schrödinger operators ${\mathsf{A}}_{\varepsilon}$ converges to the Robin Laplacian ${\mathsf{A}}$ in strong resolvent sense as $\varepsilon\to 0$.
@@
1.3 Remark (appearence of mean curvature terms in related problems).
@@
Our second main result concerns the case of non-negative potentials.
@@
Assume that the potential $V\in C^{\infty}_{\mathrm{c}}(\overline{{{\mathbb{R}}}}_{+})$ is non-negative. Then the family of operators ${\mathsf{A}}_{\varepsilon}$ converges to ${\mathsf{A}}_{0}$ in strong resolvent sense as $\varepsilon\to 0$.
The proof of this theorem relies on the same circle of ideas as the proof of Theorem A. In the construction of the identification operators we use the derivative of the non-resonant solution instead of the non-resonant solution itself as we do in the proof of Theorem A with the resonant solution. By analogy with the one-dimensional case, we expect that ${\mathsf{A}}_{\varepsilon}$ converge as $\varepsilon\to 0$ to ${\mathsf{A}}_{0}$ in strong resolvent sense for general non-resonant potential $V$. However, we have not been able to find a proof for this claim.
@@
The condition $V\geq 0$ in the non-resonant case implies that there is a constant $c>0$ (actually $c=\sqrt{2}$) such that $\lVert{\mathsf{R}}_{\varepsilon}^{*}v\rVert_{H^{1}(\Omega)}\leq c\lVert v\rVert$ for all $v\in L_{2}(\Omega)$ and all $\varepsilon$ small enough (see Lemma 4.6). If the latter estimate does not hold, then we can show that the family of forms $(\mathfrak{a}_{\varepsilon})_{\varepsilon}$ is not uniformly elliptic (see Lemma 4.7), a concept called “equi-elliptic” in [MNP13].
@@ 2.1. Tubular coordinates
In this subsection, we briefly recall main properties of tubular coordinates. For any $t>0$, we will use the notation $\Omega_{t}=\{x\in\Omega\colon\operatorname{dist}(x,\Sigma)<t\}\subset\Omega$. By [Lee, Theorem 5.25] there exists a sufficiently small $\delta>0$ such that the mapping
@@
where the derivatives $\partial_{j}$ for $j=1,\dots,d-1$ on $\Sigma$ correspond to the choice of local coordinate system. We also write $\partial_{t}$ for $\partial_{d}$ if we need to stress that the derivative is with respect to the $d$-th variable $t$.
@@ 3. Convergence for resonant potentials
We split the proof of Theorem A into several steps. We first define some auxiliary boundary mappings which will be used in a convenient representation for the difference of the sesquilinear forms of ${\mathsf{A}}$ and ${\mathsf{A}}_{\varepsilon}$. In this representation, we also use an identification operator defined in a second step and mapping functions from the form domain of ${\mathsf{A}}$ into the form domain of ${\mathsf{A}}_{\varepsilon}$ (the latter requires a Dirichlet boundary condition on $\partial\Omega$). The identification operator is basically multiplication with the scaled resonant solution $\psi_{0}$ of the bounded solution of the initial-value problem (1.1).
@@ 3.1. Auxiliary boundary mappings
$\Gamma_{t}$ $\colon\operatorname{dom}{\mathsf{A}}_{\varepsilon}\cap C^{\infty}(\overline{\Omega})\to L_{2}(\Sigma),$ $(\Gamma_{t}v)(s)$ :$\vcentcolon=\widetilde{v}(s,t),$ (3.1a)
$\Upsilon_{t}$ $\colon\operatorname{dom}{\mathsf{A}}\cap C^{\infty}(\overline{\Omega})\to L_{2}(\Sigma),$ $(\Upsilon_{t}u)(s)$ :$\vcentcolon=2\varphi(s,t)(\partial_{t}\widetilde{u})(s,t)+\partial_{t}\varphi(s,t)\widetilde{u}(s,t),$ (3.1b)
@@
for $u\in\operatorname{dom}{\mathsf{A}}\cap C^{\infty}(\overline{\Omega})$ and $v\in\operatorname{dom}{\mathsf{A}}_{\varepsilon}\cap C^{\infty}(\overline{\Omega})$ is employed. Note that the functions $\widetilde{u}$ and $\widetilde{v}$ are smooth ($\widetilde{u},\widetilde{v}\in C^{\infty}(\Sigma\times[0,\delta))$). Moreover, we have $\widetilde{u}(s,0)=0$ for all $s\in\Sigma$.
@@
$\lVert\Upsilon_{t}u\rVert_{L_{2}(\Sigma)}\leq c\sqrt{t}\lVert u\rVert_{H^{2}(\Omega_{t})}.$
@@ 3.2. The identification operator and an expression for the form difference
where $\psi_{0}\in C^{\infty}(\overline{{{\mathbb{R}}}}_{+})$ is the bounded solution of the initial-value problem (1.1) satisfying $\psi_{0}(x)=1$ for all $x>a$.
@@
[Proof]Under the assumption $\varepsilon<a^{-1}\delta$ the tubular coordinates (2.1) can be used. In particular, using the notation (3.2), we have $\widetilde{u},\widetilde{v}\in L_{2}(\Sigma\times(0,\delta);\varphi(s,t)\operatorname{d\!}\sigma(s)\operatorname{d\!}t)$ and these functions are $C^{\infty}$-smooth.
@@
where the boundary term at $t=0$ vanishes due to $\widetilde{v}(s,0)=0$ while the boundary terms at $t=a\varepsilon$ vanishes due to $\psi_{0}^{\prime}(a)=0$. Using that $\psi_{0}$ satisfies the differential equation $-\psi_{0}^{\prime\prime}+V\psi_{0}=0$ we note that the first and the last terms on the right hand side in the above formula cancel each other. The remaining two integrals just give the desired expression involving $\Gamma_{t}$ and $\Upsilon_{t}$.
@@
$\left\lvert\mathfrak{a}[u,{\mathsf{J}}_{\varepsilon}v]-\mathfrak{a}_{\varepsilon}[{\mathsf{J}}_{\varepsilon}u,v]\right\rvert$ $\leq\frac{\lVert\psi_{0}^{\prime}\rVert_{L_{\infty}}}{\varepsilon}\int_{0}^{a\varepsilon}\lVert\Upsilon_{t}u\rVert_{L_{2}(\Sigma)}\lVert\Gamma_{t}v\rVert_{L_{2}(\Sigma)}\operatorname{d\!}t$
$\leq\frac{c\lVert\psi_{0}^{\prime}\rVert_{L_{\infty}}\sqrt{a}}{\sqrt{\varepsilon}}\lVert u\rVert_{H^{2}(\Omega_{a\varepsilon})}\int_{0}^{a\varepsilon}\lVert\Gamma_{t}v\rVert_{L_{2}(\Sigma)}\operatorname{d\!}t$
$\leq ca\lVert\psi_{0}^{\prime}\rVert_{L_{\infty}}\lVert u\rVert_{H^{2}(\Omega_{a\varepsilon})}\Bigl{(}\int_{0}^{a\varepsilon}\int_{\Sigma}\lvert\widetilde{v}(s,t)\rvert^{2}\operatorname{d\!}\sigma(s)\operatorname{d\!}t\Bigr{)}^{1/2}$
$\leq\hat{c}\lVert u\rVert_{H^{2}(\Omega_{a\varepsilon})}\lVert v\rVert_{L_{2}(\Omega_{a\varepsilon})}\qquad\text{where}\qquad\hat{c}\vcentcolon=\frac{ca\lVert\psi_{0}^{\prime}\rVert_{L_{\infty}}}{\sqrt{\rho}},$ (3.5)

### v2

@@ Abstract.
We consider Schrödinger operators on a bounded, smooth domain of dimension $d\geq 2$ with Dirichlet boundary conditions and a properly scaled potential, which depends only on the distance to the boundary of the domain. Our aim is to analyse the convergence of these operators as the scaling parameter tends to zero. If the scaled potential is resonant, the limit in strong resolvent sense is a Robin Laplacian with boundary coefficient expressed in terms of the mean curvature of the boundary. A counterexample shows that norm resolvent convergence cannot hold in general in this setting. If the scaled potential is non-resonant and satisfies an explicit assumption on the smallness of the negative part, the limit in strong resolvent sense is the Dirichlet Laplacian. We conjecture that we can drop this additional assumption in the non-resonant case.
Key words and phrases:
Schrödinger operator, resonant potential, strong resolvent convergence, identification operators, mean curvature
1991 Mathematics Subject Classification:
35J10, 81Q10
@@ 1.1. Background and motivation
A typical example, where the limit depends on whether the scaled potential is resonant, is the approximation of a Schrödinger operator with a point $\delta$-interaction in three dimensions by a family of Schrödinger operators with scaled regular potentials. First partial results on this approximation are obtained in [AFH79, F72, N77]. To the best of our knowledge approximations of $\delta$-interactions were first addressed in full detail by Albeverio and Høegh-Krohn in [AH81], who proved the convergence in strong resolvent sense. The limit has a non-trivial point interaction if the scaled potential is resonant, otherwise the limit is the free Laplacian. Shortly thereafter, it was shown in [AGH82] that under rather non-restrictive additional assumptions norm instead of strong resolvent convergence holds in this approximation. By restricting to the class of radially symmetric potentials one obtains upon separation of variables a model problem on the half-line, which was considered separately by Šeba in [Š85]. He proved that a family of half-line Schrödinger operators with Dirichlet boundary conditions and locally scaled potentials converges in norm resolvent sense to the Neumann (or Robin) Laplacian on the half-line if the scaled potential is resonant and to the Dirichlet Laplacian otherwise. Further refinements and extensions of these results can be found in [DM16, SLS21].
@@
We remark that a similar phenomenon, where the limit qualitatively depends on whether the scaled potential is resonant or not, was also observed in full line Schrödinger operators with scaled potentials [ACF07, CE07] motivated by the applications to curved quantum waveguides. An extension of these results to scaled potentials with zero mean was later performed in a series of papers [GM09, GH13, G22], where the first two papers treat the one-dimensional case, while the last one deals with the two-dimensional case. In the latter papers, the scaled potentials converge to $\delta^{\prime}$ in the sense of distributions. Thus, the resulting limit may then be interpreted as a rigorous definition of a Schrödinger operator with a $\delta^{\prime}$-potential.
@@ 1.2. Resonant potentials in one dimension
Figure $1.$ The solutions $\psi_{0}$ of (1.1) for resonant and non-resonant potentials on the left and on the right, respectively.
@@
Our aim in the present paper is to address a multi-dimensional counterpart of these convergence results, which we will describe below in detail. We remark that the result of Šeba in [Š85] is more general than we stated here. He has also shown how to approximate the one-dimensional Laplacian on the half-line with Robin boundary conditions by means of replacing $1/\varepsilon^{2}$ by $(1+\beta\varepsilon)/\varepsilon^{2}$ in the operator family ${\mathsf{H}}_{\varepsilon}$ [^1] We remark that in the final formula in [Š85] for the boundary parameter in the limiting operator the parameter $\beta$ appears in the denominator and formally the case $\beta=0$ is excluded, but careful inspection of the proof shows that it applies also to the case $\beta=0$ leading to Neumann boundary conditions in the limit.. We will not address this more general case in our analysis of the multi-dimensional problem, as we already obtain a Robin-type boundary conditions by geometry.
@@ 1.3. Main results
is called the shape operator; here $T_{s}\Sigma$ is the tangent space for $\Sigma$ at $s\in\Sigma$. The eigenvalues $\kappa_{1}(s),\kappa_{2}(s),\dots,\kappa_{d-1}(s)\colon\partial\Omega\to{\mathbb{R}}$ of ${\mathsf{L}}_{s}$ are called the principal curvatures of $\partial\Omega$. In our sign convention for the shape operator, all the principal curvatures are non-negative if and only if $\Omega$ is convex. The mean curvature of $\Sigma$ at $s\in\Sigma$ is defined as usual by
@@
We adopt the notation $H^{k}(\Omega)$ for the $L_{2}$-based Sobolev space on $\Omega$ of order $k\in{\mathbb{N}}$ and $H^{s}(\partial\Omega)$ for the $L_{2}$-based Sobolev space on the boundary $\partial\Omega$ of order $s\in{\mathbb{R}}_{+}$. We denote by $\ring{H}^{k}(\Omega)$ the closure of $C^{\infty}_{\rm c}(\Omega)$ in the Sobolev space $H^{k}(\Omega)$ with respect to its norm. For a function $u\in H^{2}(\Omega)$, we use the notation $u{\restriction}_{\Sigma}\in H^{3/2}(\Sigma)$ for its trace on the boundary and $\partial_{\nu}u{\restriction}_{\Sigma}\in H^{1/2}(\Sigma)$ for the normal derivative corresponding to the normal pointing outwards of $\Omega$. The role of the one-dimensional Neumann Laplacian $\mathsf{H}$} from the previous subsection is played now by the self-adjoint Robin Laplacian in $L_{2}(\Omega)$
@@
where the Robin coefficient is expressed in terms of the mean curvature. At the points where the mean curvature vanishes, we recover locally Neumann boundary conditions. The role of the family of one-dimensional Schrödinger operators ${\mathsf{H}}_{\varepsilon}$ is played by the self-adjoint Schrödinger operator in $L_{2}(\Omega)$ defined for $\varepsilon>0$ by
@@
Assume that $V\in C^{\infty}_{\mathrm{c}}(\overline{{{\mathbb{R}}}}_{+})$ is resonant in the sense of Definition 1.1. Then the family of scaled Schrödinger operators ${\mathsf{A}}_{\varepsilon}$ converges to the Robin Laplacian $\mathsf{A}$} in strong resolvent sense as $\varepsilon\to 0$.
@@
1.3 Remark (appearance of mean curvature terms in related problems).
@@
Our second main result concerns the case of non-resonant potentials. Let us introduce the parameter
$C_{\Omega}:=\inf_{u\in\ring{H}^{1}(\Omega)\setminus\{0\}}\frac{\displaystyle\int_{\Omega}|\nabla u|^{2}\operatorname{d\!}x}{\displaystyle\int_{\Omega}|u(x)|^{2}(\operatorname{dist}(x,\Sigma))^{-2}\operatorname{d\!}x}.$ (1.7)
By [BEL, Theorem 3.2.1, discussion below Eq. (3.2.4)] we have $C_{\Omega}=\frac{1}{4}$ for bounded smooth convex domains, while for general bounded $C^{\infty}$-smooth domains $C_{\Omega}\in(0,\frac{1}{4}]$.
@@
Assume that $V\in C^{\infty}_{\mathrm{c}}(\overline{{{\mathbb{R}}}}_{+})$ is such that $\max_{t\in{\mathbb{R}}_{+}}(-t^{2}V(t))<C_{\Omega}$. Then the family of operators ${\mathsf{A}}_{\varepsilon}$ converges to ${\mathsf{A}}_{0}$ in strong resolvent sense as $\varepsilon\to 0$.
It follows from the Hardy inequality on an interval [BEL, Eq. 3.3.4] that any potential satisfying the assumption in Theorem B is non-resonant, since the Schrödinger operator in Remark 1.2 (a) is strictly positive and thus can not have eigenvalue zero. Any non-negative potential satisfies the assumption in the above theorem. Moreover, a small negative part is allowed. The assumption imposes no restriction on the positive part of the potential. The proof of this theorem relies on the same circle of ideas as the proof of Theorem A. In the construction of the identification operators we use the derivative of the non-resonant solution instead of the non-resonant solution itself as we do in the proof of Theorem A with the resonant solution. By analogy with the one-dimensional case, we expect that ${\mathsf{A}}_{\varepsilon}$ converge as $\varepsilon\to 0$ to ${\mathsf{A}}_{0}$ in strong resolvent sense for general non-resonant potential $V$. However, we have not been able to find a proof for this claim.
@@
The condition on $V$ in the non-resonant case implies that there is a constant $c>0$ such that $\lVert({\mathsf{A}}_{\varepsilon}+{\mathsf{i}})^{-1}v\rVert_{H^{1}(\Omega)}\leq c\|v\|_{L_{2}(\Omega)}$ for all $v\in L_{2}(\Omega)$ and all $\varepsilon$ small enough (see Lemma 4.6). If the latter estimate does not hold, then we can show that the family of sesquilinear forms associated with the operators ${\mathsf{A}}_{\varepsilon}$ is not uniformly elliptic (see Lemma 4.7), a concept called “equi-elliptic” in [MNP13].
@@ 2.1. Tubular coordinates
In this subsection, we briefly recall main properties of tubular coordinates. For any $t>0$, we will use the notation $\Omega_{t}=\{x\in\Omega\colon\operatorname{dist}(x,\Sigma)<t\}\subset\Omega$, where $\operatorname{dist}(\cdot,\Sigma)$ is the Euclidean distance to $\Sigma$. By [Lee, Theorem 5.25] there exists a sufficiently small $\delta>0$ such that the mapping
@@
where the derivatives $\partial_{j}$ for $j=1,\dots,d-1$ on $\Sigma$ correspond to the choice of local coordinate system and where $(G^{jk})_{j,k=1}^{d}$ stands for the inverse to the matrix $(G_{jk})_{j,k=1}^{d}$. We also write $\partial_{t}$ for $\partial_{d}$ if we need to stress that the derivative is with respect to the $d$-th variable $t$.
@@ 3. Convergence for resonant potentials
We split the proof of Theorem A into several steps. We first define some auxiliary boundary mappings which will be used in a convenient representation for the difference of the sesquilinear forms of $\mathsf{A}$} and ${\mathsf{A}}_{\varepsilon}$. In this representation, we also use an identification operator defined in a second step and mapping functions from the form domain of $\mathsf{A}$} into the form domain of ${\mathsf{A}}_{\varepsilon}$ (the latter requires a Dirichlet boundary condition on $\partial\Omega$). The identification operator is basically multiplication with the scaled resonant solution $\psi_{0}$ of the initial-value problem (1.1).
@@ 3.1. Auxiliary boundary mappings
$\Gamma_{t}$ $\colon\operatorname{dom}{\mathsf{A}}_{\varepsilon}\cap C^{\infty}(\overline{\Omega})\to L_{2}(\Sigma),$ $(\Gamma_{t}v)(s)$ $\vcentcolon=\widetilde{v}(s,t),$ (3.1a)
$\Upsilon_{t}$ $\colon\operatorname{dom}{\mathsf{A}}\cap C^{\infty}(\overline{\Omega})\to L_{2}(\Sigma),$ $(\Upsilon_{t}u)(s)$ $\vcentcolon=2\varphi(s,t)(\partial_{t}\widetilde{u})(s,t)+\partial_{t}\varphi(s,t)\widetilde{u}(s,t),$ (3.1b)
@@
for $u\in\operatorname{dom}{\mathsf{A}}\cap C^{\infty}(\overline{\Omega})$ and $v\in\operatorname{dom}{\mathsf{A}}_{\varepsilon}\cap C^{\infty}(\overline{\Omega})$ is employed. Note that the functions $\widetilde{u}$ and $\widetilde{v}$ are smooth ($\widetilde{u},\widetilde{v}\in C^{\infty}(\Sigma\times[0,\delta))$). Moreover, we have $\widetilde{v}(s,0)=0$ for all $s\in\Sigma$.
@@
$\lVert\Upsilon_{t}u\rVert_{L_{2}(\Sigma)}\leq c\sqrt{t}\lVert u\rVert_{H^{2}(\Omega_{t})}$
@@ 3.2. The identification operator and an expression for the form difference
where $\psi_{0}\in C^{\infty}(\overline{{{\mathbb{R}}}}_{+})$ is the bounded solution of the initial-value problem (1.1) satisfying $\psi_{0}(t)=1$ for all $t>a$.
@@
[Proof]Recall that $\delta>0$ is chosen as in Subsection 2.1 so that the mapping $\Phi$ in (2.1) is a diffeomorphism from $\Sigma\times(0,\delta)$ onto $\delta$-neighbourhood $\Omega_{\delta}$ of the boundary $\partial\Omega$. Under the assumption $\varepsilon<a^{-1}\delta$ the tubular coordinates (2.1) can be used. In particular, using the notation (3.2), we have $\widetilde{u},\widetilde{v}\in L_{2}(\Sigma\times(0,\delta);\varphi(s,t)\operatorname{d\!}\sigma(s)\operatorname{d\!}t)$ and these functions are $C^{\infty}$-smooth.
@@
where the boundary term at $t=0$ vanishes due to $\widetilde{v}(s,0)=0$ while the boundary term at $t=a\varepsilon$ vanishes due to $\psi_{0}^{\prime}(a)=0$. Using that $\psi_{0}$ satisfies the differential equation $\psi_{0}^{\prime\prime}+V\psi_{0}=0$ we note that the first and the last terms on the right hand side in the above formula cancel each other. The remaining two integrals just give the desired expression involving $\Gamma_{t}$ and $\Upsilon_{t}$.
@@
$\left\lvert\mathfrak{a}[u,{\mathsf{J}}_{\varepsilon}v]-\mathfrak{a}_{\varepsilon}[{\mathsf{J}}_{\varepsilon}u,v]\right\rvert$ $\leq\frac{\lVert\psi_{0}^{\prime}\rVert_{\infty}}{\varepsilon}\int_{0}^{a\varepsilon}\lVert\Upsilon_{t}u\rVert_{L_{2}(\Sigma)}\lVert\Gamma_{t}v\rVert_{L_{2}(\Sigma)}\operatorname{d\!}t$
$\leq\frac{c\lVert\psi_{0}^{\prime}\rVert_{\infty}\sqrt{a}}{\sqrt{\varepsilon}}\lVert u\rVert_{H^{2}(\Omega_{a\varepsilon})}\int_{0}^{a\varepsilon}\lVert\Gamma_{t}v\rVert_{L_{2}(\Sigma)}\operatorname{d\!}t$
$\leq ca\lVert\psi_{0}^{\prime}\rVert_{\infty}\lVert u\rVert_{H^{2}(\Omega_{a\varepsilon})}\Bigl(\int_{0}^{a\varepsilon}\int_{\Sigma}\lvert\widetilde{v}(s,t)\rvert^{2}\operatorname{d\!}\sigma(s)\operatorname{d\!}t\Bigr)^{1/2}$
$\leq\hat{c}\lVert u\rVert_{H^{2}(\Omega_{a\varepsilon})}\lVert v\rVert_{L_{2}(\Omega_{a\varepsilon})}\qquad\text{where}\qquad\hat{c}\vcentcolon=\frac{ca\lVert\psi_{0}^{\prime}\rVert_{\infty}}{\sqrt{\rho}},$ (3.5)